OpenStudy (anonymous):

HE lp !! !! Find the 9th term of the sequence described by: A(n) = 7 + (n-1)(3) A) 42 B) 9 C) 31 D) 27

OpenStudy (anonymous):

where you see an \(n\) put a \(9\) compute \[A_9=7+(9-1)\times 3\] that is all

OpenStudy (anonymous):

Do u mean A^9 ?

OpenStudy (alexandervonhumboldt2):

a(9)=7+24=31

OpenStudy (alexandervonhumboldt2):

so the 9th term is 31

OpenStudy (anonymous):

so C ?

OpenStudy (anonymous):

at the risk of repeating myself, replace \(n\) by \(9\)!

OpenStudy (anonymous):

\[A_9=7+(9-1)\times 3\] see what you get when you do that piece of arithmetic

OpenStudy (anonymous):

yeah but do u mean A^9

OpenStudy (anonymous):

does it say \(A^9\) ?

OpenStudy (anonymous):

no but i don't know how to do the other one

OpenStudy (anonymous):

i wish i knew another way to put it you want \(A_{\color{red}9}\) and you have \[A_\color{red}n=7+(\color{red}n-1)\times 3\] therefore \[\huge A_{\color{red}9}=7+(\color{red}9-1)\times 3\]

OpenStudy (anonymous):

does A 9 mean Ax9 ?

OpenStudy (anonymous):

Dan, help me out here

OpenStudy (anonymous):

no it means the ninth term

OpenStudy (anonymous):

the first term is \(A_1\) the second is \(A_2\) the third is \(A_3\) and the nth is \(A_n\) you want the ninth term

OpenStudy (anonymous):

C ?

OpenStudy (anonymous):

31 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

THX !

OpenStudy (anonymous):

yw